OpenStudy (smarty):

Solve 3x^2 + 6x + 6 = 0 using the quadratic equation. A.x = 1/2, x = –3 B.x = –1 + i, x = –1 – i C.x = –2, x = 2 D.x = 6 + i, x = 6 – i

OpenStudy (smarty):

I got: And i don't know what to do afterwards. 

OpenStudy (anonymous):

you have the right answer i believe

OpenStudy (anonymous):

simplify it by splittng the denominator up with the two numbers

OpenStudy (smarty):

So would my answer be D?

Parth (parthkohli):

\(\sqrt{36} = \sqrt{-1} \times \sqrt{36}= 6i\) You were going in the right direction :)

Parth (parthkohli):

No

OpenStudy (anonymous):

B

OpenStudy (anonymous):

it's b not d

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {6(1 \pm i) \over 6}}\) \( \color{Black}{\Rightarrow 1 \pm i}\)

Parth (parthkohli):

Get it?